# Foldable SuperNets: Scalable Merging of Transformers with Different Initializations and Tasks

### v1

@@ B.1.3 The Multi-Layer Perceptron block
The Foldable SuperNet of the $l$ MLP block receives $Y_{l}^{*}\in\mathbb{R}^{T\times d}$ as input, which are the merged features of the previous attention block. It learns the layer norm parameters of the MLP block ${\color[rgb]{1,0,0}\alpha^{*}_{l},\theta^{*}_{l}}\in\mathbb{R}^{d}$, which, as usual, acts on the compressed dimension:
$\bar{Y}_{l}^{*}=\text{LN}_{{\color[rgb]{1,0,0}\alpha^{*}_{l},\theta^{*}_{l}}}(Y_{l}^{*})\,.$
@@
$X_{l+1}^{*}=Y_{l}^{*}+\sigma\left(\bar{Y}_{l}^{*}{\color[rgb]{1,0,0}U_{l}^{1}}\begin{pmatrix}W_{l}^{1,A}&0\\ 0&W_{l}^{1,B}\end{pmatrix}{\color[rgb]{1,0,0}M_{l}^{1}}\right){\color[rgb]{1,0,0}U_{l}^{2}}\begin{pmatrix}W_{l}^{2,A}&0\\ 0&W_{l}^{2,B}\end{pmatrix}{\color[rgb]{1,0,0}M_{l}^{2}}\,.$
Where ${\color[rgb]{1,0,0}U_{l}^{1},\,U_{l}^{2}}\in\mathbb{R}^{d\times 2\cdot d}$, ${\color[rgb]{1,0,0}M_{l}^{1},\,M_{l}^{2}}\in\mathbb{R}^{2\cdot d\times d}$. $X_{l+1}^{*}\in\mathbb{R}^{T\times d}$ then serves as the input for the $l+1$ attention block.
@@ B.2 Training the Foldable SuperNet
In the case of merging ViTs, this global approach worked much better than addressing a series of local problems for each block, as was done in the MLP case. For more details, please refer to Appendix H.1.
@@ Appendix C Time and memory complexity analysis
FS-Merge low rank $\sim$ 3.6 Hours 60M 91.54
FS-Merge seq. $\sim$ 2.2 Hours 18M 90.94
@@ D.1 Training Vision Transformers from scratch
We pre-trained ViT-B-16 and ViT-L-14 (dosovitskiy2021an; steiner2022how; touvron2021training) on ImageNet-1K (deng2009imagenet). These models were initialized from distinct random seeds and exposed to training data in different orders. Following a setting similar to other merging works (ilharco2023editing; stoica2024zipit), a frozen classification head derived from CLIP’s (radford2021learning) label embeddings was used, in order to make the outputs space of the ViTs similar. Training and merging ViTs with learned classification heads are left for future research. It is important to mention that the classification heads are not being merged.
We pre-trained the ViTs following common practices (touvron2021training), such as augmentations, MixUp (zhang2018mixup), distillation using resnet152 (he2016deep), cross-entropy loss, and a cosine scheduler with a single cycle and a warmup.
@@ D.2 Datasets and fine-tuned Vision Transformers
After obtaining different pre-trained ViTs, we fine-tuned each ViT on a different downstream task. Following the approach outlined in (ilharco2023editing), we fine-tuned on a range of datasets including Cars (krause20133d), DTD (cimpoi2014describing), EuroSAT (helber2019eurosat), GTSRB (stallkamp2011german), MNIST (lecun1998mnist), RESISC45 (cheng2017remote), SVHN (netzer2011reading), CIFAR10, and CIFAR100 (krizhevsky2009learning). We used the existing datasets’ training set, validation set, and test set. If there wasn’t a validation set, one was created by using 15% of the training set.
All models were fine-tuned with a batch size of 256, a learning rate of $1e^{-5}$, cross-entropy loss, and a cosine scheduler using a single cycle with a warm-up phase. As done in previous works (ilharco2023editing; stoica2024zipit), the classification head was frozen and obtained from CLIP embeddings (radford2021learning) of the label names. In Table 7, we present the dataset details and the test accuracy of the fine-tuned models on their respective tasks. The models’ accuracies on tasks they were not fine-tuned for are as good as a random guess.
@@ E.1 Merging Multi-Layer Perceptrons
We divided the MNIST dataset (lecun1998mnist) into two subsets: the first containing labels 0 to 4 and the second containing labels 5 to 9. Each subset was further split into training, validation, and testing sets, with the validation set comprising 10% of the original training set. A similar approach was adopted for the Fashion-MNIST dataset (xiao2017fashion).
We trained MLPs on each of the MNIST subsets, utilizing SGD with a learning rate of 0.01, a batch size of 10, and cross-entropy loss. The training duration was set to 1 epoch. However, if the number of hidden layers exceeded four, or the hidden dimension surpassed 256, the training was extended to 2 epochs. For the Fashion MNIST dataset (xiao2017fashion), the same hyperparameters were used, with adjustments made only to the number of epochs, increased to 10, and the learning rate, reduced to 0.001.
@@ E.2 Merging Vision Transformers
Data and augmentations. We took a set of images from the training set and expanded it using augmentations. Specifically, these augmentations were used: Random Crop; Random Horizontal Flip; Random choice between grayscale, Solarization, and GaussianBlur; and MixUp (zhang2018mixup). Then this dataset is used to create features, where features of the model fine-tuned on task $A$ were generated only from images of this task (and their augmentations). For efficiency and reproducibility, we saved these features and then used them for all our merging methods. In the case of FS-Merge and distillation, a single epoch means using the entire features dataset once, including the features created by augmentations.
@@
RegMean (jin2023dataless). $\alpha$ is a factor which decrease the non-diagonal items of the inner product matrices in the RegMean solution.
@@
“Opt” (imfeld2023transformer). The hyperparameters include the filter type (which determines the feature tokens used by the optimal transport solver) and $\lambda$, a regularization term for the solver (Sinkhorn-Knapp algorithm). Lower values of $\lambda$ result in harder alignment.
@@ F.1 Merging Multi-Layer Perceptrons
FS-Merge. Our method was tested in five variants. FS-Merge is the local version (Eq. 3), which trains a Foldable SuperNet for each layer $l$ independently, using the original models’ pre-activation features as inputs $z_{l}^{A},z_{l}^{B}$. FS-Merge global is the global version of our method (Eq. 4), training a Foldable SuperNet for all the layers together to reconstruct the features of the final representation layer $f_{L}^{A},f_{L}^{B}$. For both of these versions, we also tested a variant where we initialized the Foldable SuperNet’s $M$ and $U$ with the solutions of ZipIt (stoica2024zipit), and then continued to optimize them (FS-Merge ZipIt init and FS-Merge global ZipIt init). FS-Merge no cross compresses each of the two MLPs ($A$ and $B$) individually to half of their widths using a local FS-Merge and then concatenates those two compressed models. This means that neurons between these two models cannot be merged.
@@
Table 9 presents the per-task accuracy for merging MLPs trained on half of the MNIST dataset (lecun1998mnist), with varying hidden width. Table 10 and Table 11 present the joint accuracy for fusing MLPs trained on half of the MNIST dataset, with variations in depth or hidden width, respectively. Similarly, Table 12 and Table 13 present the per-task accuracy for merging MLPs trained on half of the Fashion MNIST dataset (xiao2017fashion), again varying by depth or hidden width. The joint accuracy for these models are detailed in Table 14.
@@ F.2 Merging Vision Transformers
Baselines. Our goal was to merge Vision Transformers (ViTs) that were initialized differently and trained on various tasks. We compared our method against “average” (wortsman2022model), a simple weight averaging technique; “SLERP” (shoemake1985animating), spherical linear interpolation; RegMean (jin2023dataless), which offers a closed-form solution by solving a linear regression problem for each linear layer; “Opt” (imfeld2023transformer), which uses optimal transport (knight2008sinkhorn) to align transformers, and can be viewed as a generalization of neuron alignment methods (ainsworth2023git) because it can find soft alignments as well as hard ones; and a distillation baseline (hinton2015distilling), which trains a single ViT to mimic the features of the last representation layer of the original models.
@@
Table 20 and Table 21 present additional results from the experiment involving the merging of ViT-L-14 pairs. This experiment investigated how the merging methods scale to larger models, with FS-Merge Low Rank applied using a rank of 32.
@@
FS-Merge, low rank 72.93 39.48 85.69 82.10 91.54 77.63
FS-Merge seq. 74.11 39.63 84.64 75.93 90.66 79.68
@@
FS-Merge, low rank 69.45 60.48 67.44 36.81 67.78 59.80 60M
FS-Merge seq. 70.05 58.08 67.23 38.76 65.26 55.20 18M
@@
FS-Merge, low rank 78.20 65.90 76.13 66.91 79.71 64.94
FS-Merge seq. 76.05 50.64 72.93 61.55 78.28 60.65
@@
FS-Merge diagonal 85.93 77.27 63.79 61.36 91.99 84.8
FS-Merge low rank 88.10 79.33 69.23 67.22 92.41 89.74
F.3 Merging text Transformers

### v2

@@ B.1.3 The Multi-Layer Perceptron block
The Foldable SuperNet of the $l$ MLP block receives $Y_{l}^{*}\in\mathbb{R}^{T\times d}$ as input, which are the merged features of the previous attention block. It learns the layer norm parameters of the MLP block ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha^{*}_{l},\theta^{*}_{l}}\in\mathbb{R}^{d}$, which, as usual, acts on the compressed dimension:
$\bar{Y}_{l}^{*}=\text{LN}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha^{*}_{l},\theta^{*}_{l}}}(Y_{l}^{*})\,.$
@@
$X_{l+1}^{*}=Y_{l}^{*}+\sigma\left(\bar{Y}_{l}^{*}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}^{1}}\begin{pmatrix}W_{l}^{1,A}&0\\ 0&W_{l}^{1,B}\end{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{1}}\right){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}^{2}}\begin{pmatrix}W_{l}^{2,A}&0\\ 0&W_{l}^{2,B}\end{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{2}}\,.$
Where ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}^{1},\,U_{l}^{2}}\in\mathbb{R}^{d\times 2\cdot d}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{1},\,M_{l}^{2}}\in\mathbb{R}^{2\cdot d\times d}$. $X_{l+1}^{*}\in\mathbb{R}^{T\times d}$ then serves as the input for the $l+1$ attention block.
@@ B.2 Training the Foldable SuperNet
In the case of merging ViTs, this global approach worked much better than addressing a series of local problems for each block, as was done in the MLP case. For more details, please refer to Appendix G.3.
@@ Appendix C Time and memory complexity analysis
FS-Merge low rank $\sim$ 3.6 Hours 60M 91.63
FS-Merge seq. $\sim$ 2.2 Hours 18M 91.10
@@ D.1 Training Vision Transformers from scratch
We pre-trained ViT-B-16 and ViT-L-14 (Dosovitskiy et al., 2021; Steiner et al., 2022; Touvron et al., 2021) on ImageNet-1K (Deng et al., 2009). These models were initialized from distinct random seeds and exposed to training data in different orders. Following a setting similar to other merging works (Ilharco et al., 2023; Stoica et al., 2024), a frozen classification head derived from CLIP’s (Radford et al., 2021) label embeddings was used, in order to make the outputs space of the ViTs similar. Training and merging ViTs with learned classification heads are left for future research. It is important to mention that the classification heads are not being merged.
We pre-trained the ViTs following common practices (Touvron et al., 2021), such as augmentations, MixUp (Zhang et al., 2018), distillation using resnet152 (He et al., 2016), cross-entropy loss, and a cosine scheduler with a single cycle and a warmup.
@@ D.2 Datasets and fine-tuned Vision Transformers
After obtaining different pre-trained ViTs, we fine-tuned each ViT on a different downstream task. Following the approach outlined in (Ilharco et al., 2023), we fine-tuned on a range of datasets including Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), RESISC45 (Cheng et al., 2017), SVHN (Netzer et al., 2011), CIFAR10, and CIFAR100 (Krizhevsky et al., 2009). We used the existing datasets’ training set, validation set, and test set. If there wasn’t a validation set, one was created by using 15% of the training set.
All models were fine-tuned with a batch size of 256, a learning rate of $1e^{-5}$, cross-entropy loss, and a cosine scheduler using a single cycle with a warm-up phase. As done in previous works (Ilharco et al., 2023; Stoica et al., 2024), the classification head was frozen and obtained from CLIP embeddings (Radford et al., 2021) of the label names. In Table 8, we present the dataset details and the test accuracy of the fine-tuned models on their respective tasks. The models’ accuracies on tasks they were not fine-tuned for are as good as a random guess.
@@ E.1 Merging Multi-Layer Perceptrons
We divided the MNIST dataset (LeCun, 1998) into two subsets: the first containing labels 0 to 4 and the second containing labels 5 to 9. Each subset was further split into training, validation, and testing sets, with the validation set comprising 10% of the original training set. A similar approach was adopted for the Fashion-MNIST dataset (Xiao et al., 2017).
We trained MLPs on each of the MNIST subsets, utilizing SGD with a learning rate of 0.01, a batch size of 10, and cross-entropy loss. The training duration was set to 1 epoch. However, if the number of hidden layers exceeded four, or the hidden dimension surpassed 256, the training was extended to 2 epochs. For the Fashion MNIST dataset (Xiao et al., 2017), the same hyperparameters were used, with adjustments made only to the number of epochs, increased to 10, and the learning rate, reduced to 0.001.
@@ E.2 Merging Vision Transformers
Data and augmentations. We took a set of images from the training set and expanded it using augmentations. Specifically, these augmentations were used: Random Crop; Random Horizontal Flip; Random choice between grayscale, Solarization, and GaussianBlur; and MixUp (Zhang et al., 2018). Then this dataset is used to create features, where features of the model fine-tuned on task $A$ were generated only from images of this task (and their augmentations). For efficiency and reproducibility, we saved these features and then used them for all our merging methods. In the case of FS-Merge and distillation, a single epoch means using the entire features dataset once, including the features created by augmentations.
@@
RegMean (Jin et al., 2023). $\alpha$ is a factor which decrease the non-diagonal items of the inner product matrices in the RegMean solution.
@@
“Opt” (Imfeld et al., 2023). The hyperparameters include the filter type (which determines the feature tokens used by the optimal transport solver) and $\lambda$, a regularization term for the solver (Sinkhorn-Knapp algorithm). Lower values of $\lambda$ result in harder alignment.
@@ F.1 Merging Multi-Layer Perceptrons
FS-Merge. Our method was tested in five variants. FS-Merge is the local version (Eq. 3), which trains a Foldable SuperNet for each layer $l$ independently, using the original models’ pre-activation features as inputs $z_{l}^{A},z_{l}^{B}$. FS-Merge global is the global version of our method (Eq. 4), training a Foldable SuperNet for all the layers together to reconstruct the features of the final representation layer $f_{L}^{A},f_{L}^{B}$. For both of these versions, we also tested a variant where we initialized the Foldable SuperNet’s $M$ and $U$ with the solutions of ZipIt (Stoica et al., 2024), and then continued to optimize them (FS-Merge ZipIt init and FS-Merge global ZipIt init). FS-Merge no cross compresses each of the two MLPs ($A$ and $B$) individually to half of their widths using a local FS-Merge and then concatenates those two compressed models. This means that neurons between these two models cannot be merged.
@@
Table 10 presents the per-task accuracy for merging MLPs trained on half of the MNIST dataset (LeCun, 1998), with varying hidden width. Table 11 and Table 12 present the joint accuracy for fusing MLPs trained on half of the MNIST dataset, with variations in depth or hidden width, respectively. Similarly, Table 13 and Table 14 present the per-task accuracy for merging MLPs trained on half of the Fashion MNIST dataset (Xiao et al., 2017), again varying by depth or hidden width. The joint accuracy for these models are detailed in Table 15.
@@ F.2 Merging Vision Transformers
Baselines. Our goal was to merge Vision Transformers (ViTs) that were initialized differently and trained on various tasks. We compared our method against “average” (Wortsman et al., 2022), a simple weight averaging technique; “SLERP” (Shoemake, 1985), spherical linear interpolation; RegMean (Jin et al., 2023), which offers a closed-form solution by solving a linear regression problem for each linear layer; “Opt” (Imfeld et al., 2023), which uses optimal transport (Knight, 2008) to align transformers, and can be viewed as a generalization of neuron alignment methods (Ainsworth et al., 2023) because it can find soft alignments as well as hard ones; and a distillation baseline (Hinton et al., 2015), which trains a single ViT to mimic the features of the last representation layer of the original models.
@@
Table 21, Table 22 and Table 23 present additional results from the experiment involving the merging of ViT-L-14 pairs. This experiment investigated how the merging methods scale to larger models, with FS-Merge Low Rank applied using a rank of 32.
@@
FS-Merge, low rank 73.26 41.33 86.44 78.78 91.38 79.12
FS-Merge seq. 74.10 40.76 84.85 75.86 90.93 80.29
@@
FS-Merge, low rank 71.84 65.43 68.12 38.11 68.31 60.84 60M
FS-Merge seq. 71.71 61.21 67.75 38.14 66.19 56.64 18M
@@
FS-Merge, low rank 77.05 60.77 75.97 65.95 80.83 65.50
FS-Merge seq. 76.82 55.30 72.80 62.51 78.24 60.36
Merging Methods DTD, EuroSAT CIFAR100, SVHN Cars, MNIST
Per-task Joint Per-task Joint Per-task Joint
Original models 81.33 - 94.68 - 96.50 -
Ensemble 77.71 71.54 94.25 77.90 96.23 96.22
Average 6.36 1.40 9.18 8.03 5.98 0.08
SLERP 5.21 1.58 5.20 2.59 8.48 5.44
RegMean 8.66 4.21 5.64 0.47 10.97 0.13
Opt 10.33 3.11 4.68 2.45 5.95 5.67
Distillation 78.51 75.84 90.91 85.78 91.82 90.58
FS-M 78.60 74.68 91.78 87.67 95.80 94.31
@@
FS-Merge diagonal 85.93 77.27 63.79 61.36 91.99 84.80
FS-Merge low rank 88.11 78.32 69.09 67.14 93.56 88.20
F.3 Merging Vision Transformers with different pre-trained strategies
The next section investigates the ability of merging methods to merge vision transformers that used different pre-training strategies. ViT-B-16 models were pre-trained from scratch using supervised training on the ImageNet-1k dataset (Deng et al., 2009) (the pre-training strategy used throughout the rest of this work). These models were then fine-tuned on RESISC45 (Cheng et al., 2017), DTD (Cimpoi et al., 2014), SVHN (Netzer et al., 2011), and EuroSAT (Helber et al., 2019). Additionally, the vision encoder from a pre-trained CLIP model (Radford et al., 2021), which is also a ViT-B-16, was used. This ViT was pre-trained from scratch using a contrastive learning approach that leverages pairs of similar and dissimilar image-caption pairs. Subsequently, this ViT was fine-tuned on Cars (Krause et al., 2013), GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), CIFAR10, and CIFAR100 (Krizhevsky et al., 2009).
Table 24 presents experiments in which a CLIP pre-trained model was merged with an ImageNet pre-trained model, using the CLIP model as the “first” initialization. Table 25 shows similar experiments, using the ImageNet pre-trained model as the “first” initialization. Note that the order of the models is important due to the “first” initialization used by both FS-Merge and distillation. For merging, 16 original images from the training set and 800 augmented images were used from each dataset (a total of 1,632 images per merge). The merged model was then evaluated on the fine-tuned datasets. FS-Merge was used with low rank of 12.
As shown, even with the new pre-training strategy, our main conclusions from the other experiments remain consistent: local and simple merging methods, such as Average and Opt, fail in this setting, while FS-Merge continues to outperform distillation by a significant margin in most cases.
Table 24: Merging pairs of ViT-B-16 models, where the first model is pre-trained using CLIP and the second model is pre-trained on ImageNet. The merge is performed using 16 original images from the training set and 800 augmented images from each dataset. We report the per-task and joint accuracy on the test set.
Merging Methods Cars, RESISC45 GTSRB, EuroSAT CIFAR10, DTD
Per-task Joint Per-task Joint Per-task Joint
Original models 90.08 - 98.99 - 81.26 -
Ensemble 82.40 81.94 97.72 95.18 80.58 79.88
Average 1.76 1.49 6.39 2.74 6.26 4.95
SLERP 1.50 1.14 6.11 2.60 5.89 2.26
Opt 1.87 1.20 5.75 2.18 6.10 3.61
Distillation 51.10 46.62 63.96 61.26 56.36 55.48
FS-Merge 60.36 59.52 84.52 84.23 58.56 58.11
Table 25: Merging pairs of ViT-B-16 models, where the first model is pre-trained using ImageNet and the second model is pre-trained on CLIP. The merge is performed using 16 original images from the training set and 800 augmented images from each dataset. We report the per-task and joint accuracy on the test set.
Merging Methods RESISC45, GTSRB DTD, MNIST SVHN, CIFAR100
Per-task Joint Per-task Joint Per-task Joint
Original models 96.19 - 81.86 - 93.83 -
Ensemble 92.96 91.14 83.01 69.70 82.38 72.74
Average 2.79 1.73 3.10 1.30 8.19 0.83
SLERP 2.45 1.43 6.59 1.43 8.08 0.48
Opt 1.86 1.02 4.75 1.32 7.21 0.47
Distillation 60.08 57.63 67.62 67.35 50.25 43.12
FS-Merge 64.55 62.43 73.90 70.97 51.74 40.06
F.4 Merging Vision Transformers fine-tuned from the same initialization
To explore FS-Merge’s capabilities, we investigate its performance in the more common setting found in merging method literature: merging models fine-tuned on different tasks from the same pre-trained initialization. We fine-tuned the ViT-B-16 model on GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), RESISC45 (Cheng et al., 2017), CIFAR10, and CIFAR100 (Krizhevsky et al., 2009), all starting from the same pre-trained model. Subsequently, we merged groups of these fine-tuned models in varying sizes.
The results, shown in Table 26, utilize FS-Merge and distillation with the “average” initialization, as this configuration yielded better accuracy on the validation set. As demonstrated, FS-Merge outperforms other baselines in most cases. However, we note that in this setting, FS-Merge is less dominant, as its margin over other baselines, such as distillation and RegMean, is relatively small. Moreover, when the models are fine-tuned from the same pre-trained model, traditional merging methods, such as RegMean, achieve performance very close to FS-Merge while using significantly fewer resources. Thus, despite its success in this scenario, we conclude that FS-Merge’s true strength lies in more challenging scenarios involving models with different initializations.
Table 26: Merging ViT-B-16 models, fine-tuned on different tasks from the same pre-trained initialization. The merge is performed using 100 original images from the training set and 1000 augmented images from each dataset. The per-task and joint accuracy on the test set are reported. We will denote: C10 = CIFAR10, C100 = CIFAR100, G = GTSRB, M = MNIST, R = RESISC45.
Merging Methods G, R R, C10, G C100, R, M, G
Per-task Joint Per-task Joint Per-task Joint
Original models 96.135 - 96.68 - 94.45 -
Ensemble 92.66 86.85 86.15 67.88 80.40 41.91
Average 82.24 72.43 59.00 37.71 41.42 12.20
SLERP 82.23 72.42 57.08 45.68 38.76 30.82
RegMean 93.56 90.35 87.51 76.29 75.72 46.73
Distillation 91.76 90.37 90.85 86.44 82.41 63.11
FS-Merge 93.30 91.41 91.62 86.52 84.51 58.43
F.5 Merging Vision Transformers on large datasets
To evaluate the merging methods on larger datasets, different pre-trained models were fine-tuned on SUN397 Xiao et al. (2010), Food101 (Bossard et al., 2014), and CIFAR100 (Krizhevsky et al., 2009). Combined, these three datasets contain over 290,000 images and 598 classes. This represents the most challenging merging attempt known to us in the literature under our setting (merging differently initialized transformers with only a small unlabeled subset of the training data). Table 27 presents the results of merging these models using 100 original images and 100 augmented images per dataset. FS-Merge was applied with a low rank of 32. As shown, FS-Merge outperforms the baselines even in this challenging scenario.
Table 27: Merging ViT-B-16 models, fine-tuned on SUN397, Food101 and CIFAR100. The merge is performed using 100 original images from the training set and 1000 augmented images from each dataset. The test accuracies for individual tasks, along with the per-task and joint accuracies on the test set, are reported.
Merging Methods SUN397 Food101 CIFAR100 Per-task accuracy Joint accuracy
Original models 65.37 81.93 85.61 77.63 -
Ensemble 46.91 69.13 74.96 63.66 55.61
Average 0.10 0.95 0.50 0.51 0.05
SLERP 0.15 0.95 0.74 0.61 0.32
Opt 0.12 0.90 0.65 0.55 0.24
Distillation 44.20 20.46 35.19 33.28 27.17
FS-Merge 47.71 26.43 38.87 37.67 32.82
F.6 Merging large number of Vision Transformers
To investigate the ability of FS-Merge Seq. in the extremely challenging setting of merging a large number of models, we merged all our existing ViT-B-16 models, which were trained from different initializations on different tasks—a total of eight models. We used a relatively small amount of data: 100 original images and 1000 augmented images per task. The models were trained on Cars, DTD, CIFAR100, GTSRB, RESISC45, MNIST, EuroSAT, and SVHN, which together comprise around 5000 classes.
As shown in Table 28, this task is difficult, even for the ensemble method, which has direct access to all eight models. FS-Merge Seq. outperforms the baselines and even achieves higher joint accuracy than the ensemble, demonstrating its superiority.
However, we note that a key challenge in this scenario is the computational cost of merging such a large number of models: FS-Merge was not evaluated due to its high memory requirements for storing the frozen weights of all models. Additionally, the KD process took around 4.0 hours, while FS-Merge Seq. required approximately 5.3 hours. In contrast, traditional merging methods execute much faster: averaging and SLERP take only a few seconds, RegMean takes around 6 minutes, and Opt requires approximately 41 minutes. However, all of these methods are inadequate in this challenging setting, resulting in merged models that perform no better than random guessing.
Table 28: Merging eight ViT-B-16 models trained on different tasks using 100 original and 1000 augmented images per task. We report per-task and joint accuracy on the test set.
Metric Ensemble Average SLERP Opt RegMean Distillation FS-Merge seq.
Per-task accuracy 70.53 4.44 5.45 4.78 5.03 63.72 68.81
Joint accuracy 42.37 0.23 0.17 0.11 0.17 42.98 46.00
F.7 Merging text Transformers